find and equation of the line having the given slope and containing the given point
m=2/3, (7,-4), the equation of the line is y=?

Answer by ewatrrr(10682)


Hi


Line with m=2/3 passing thru Pt(7,-4)
y = (2/3)x + b |using ordered pair Pt(7,-4) to solve for b
-4 = 2/3*7 + b
-26/3 = b
y = (2/3)x -26/3 | Pt(0,-26/3) the y-intercept
Note: as y increases/decreases by 2, x increases/decreases by 3, thus Pt(3,-20/3)